OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (3, 2) and (5, 12)

OpenStudy (anonymous):

slope or m= (y2-y1)/(x2-x1)

OpenStudy (anonymous):

ohhh Lol

OpenStudy (anonymous):

m = (2-12)/(3-5) = -10/-2 = 5